OpenStudy (anonymous):

how large should n be to guarantee that the Simpson's Rule approximation to \[\int_{0}^{1} e^{x^2} dx\] is accurate to within 0.00001?

OpenStudy (anonymous):

Simpson's error related to h^2=k/n^2 so 1/n^2=1/100000_>n>100sq(10)#340

OpenStudy (anonymous):

\[h^2 = \frac {k}{n^2}\] \[\frac{1}{n^2}=\frac{1}{100000} \] >n>100\sq(10) #340 ?

OpenStudy (anonymous):

yes !

OpenStudy (anonymous):

means you should select n bigger than 340

OpenStudy (anonymous):

how did you come to that conclusion?

OpenStudy (anonymous):

sorry sq(10)=3.16 so answer is 317 or more.

OpenStudy (anonymous):